# Relational Prompt-based Pre-trained Language Models for Social Event Detection

### v1

@@ Abstract.
Social Event Detection (SED) aims to identify significant events from social streams, and has a wide application ranging from public opinion analysis to risk management. In recent years, Graph Neural Network (GNN) based solutions have achieved state-of-the-art performance. However, GNN-based methods often struggle with noisy and missing edges between messages, affecting the quality of learned message embedding. Moreover, these methods statically initialize node embedding before training, which, in turn, limits the ability to learn from message texts and relations simultaneously. In this paper, we approach social event detection from a new perspective based on Pre-trained Language Models (PLMs), and present $\mathrm{RPLM}_{SED}$ (Relational prompt-based Pre-trained Language Models for Social Event Detection). We first propose a new pairwise message modeling strategy to construct social messages into message pairs with multi-relational sequences. Secondly, a new multi-relational prompt-based pairwise message learning mechanism is proposed to learn more comprehensive message representation from message pairs with multi-relational prompts using PLMs. Thirdly, we design a new clustering constraint to optimize the encoding process by enhancing intra-cluster compactness and inter-cluster dispersion, making the message representation more distinguishable. We evaluate the $\mathrm{RPLM}_{SED}$ on three real-world datasets, demonstrating that the $\mathrm{RPLM}_{SED}$ model achieves state-of-the-art performance in offline, online, low-resource, and long-tail distribution scenarios for social event detection tasks.
@@ 1. Introduction
However, GNN-based social event detection methods still exhibit the following limitations. Firstly, graphs constructed from common attributes (e.g., entities, hashtags, users, etc.) between messages face issues of missing and noisy edges (Liu et al., 2022). These issues of missing and noisy edges are primarily manifested in the lack of connections between message nodes within the same event and the existence of connections between nodes that belong to different events. Whether modeling social messages as HIN graph (Figure 1(a)), homogeneous graph (Figure 1(b)), or multi-relational graph (Figure 1(c)), none of these approaches can handle well the issue of missing edges between messages of intra-event and noisy edges between massages of inter-events. This challenge is multifaceted. On the one hand, messages from the same event may have no common attributes, leading to message nodes lacking connections. This impedes GNNs from capturing associations between these nodes, resulting in an incomplete representation of related information. On the other hand, attribute co-occurrence is probable for messages from events that are not within the same event, potentially introducing noise and impacting the quality of node representation. Besides, in scenarios where the graph is sparse, GNN-based methods fail to transmit information adequately, leading to suboptimal message representations. Secondly, these GNN-based social event detection methods initialize node embeddings statically via pre-trained word embedding models (Mikolov et al., 2013; Pennington et al., 2014; Bojanowski et al., 2017) before training. This practice prevents the acquisition of better semantic embeddings from the message content in subsequent learning processes (Yang et al., 2021). Meanwhile, GNNs inherently separate the utilization of relations/edges and message content in the process of learning message embeddings from explicit message graphs. The role of edges primarily selects nodes for aggregation rather than directly participating in the processing of message representations. This approach may not fully consider the complexity of relations between messages and the potential interactions between these relations and message content (Ding et al., 2023). Thirdly, GNN-based social event detection methods exhibit insufficient constraints on the training process, treating messages merely as targets to be pulled closer or pushed apart (Hermans et al., 2017).
@@
Herein, to circumvent the aforementioned challenges of GNN-based methods, we put forward a perspective of learning expressive message representations with Pre-trained Language Models (PLMs), instead of GNNs. Thus, we present a novel Relational prompt-based Pre-trained Language Model for Social Event Detection, namely $\mathrm{RPLM}_{SED}$. Concretely, we first propose a new pairwise message modeling strategy to model social messages. As illustrated in Figure 1(d), we sample other messages for each message to form pairs and convert the multiple relations between the two into a multi-relational sequence to preserve structural information. This strategy addresses the issues of missing and noisy edges in the message graph and achieves the joint modeling of the messages’ structure and semantics. Secondly, a new multi-relational prompt-based pairwise message learning mechanism is proposed to learn more detailed and comprehensive message representations from semantic and structural information between messages. We transform the multi-relational sequences into multi-relational prompt embeddings and integrate them with the content embeddings of messages. By consistently encoding the embeddings of message pairs along with their corresponding multi-relational prompt embeddings through PLMs (Kenton and Toutanova, 2019; Liu et al., 2019; Zhang et al., 2023; Loureiro et al., 2023), we achieve the simultaneous utilization of message content and structural information, as well as the dynamic embedding and encoding of message content. To obtain more robust and stable message representations from their corresponding message pairs, a similarity-based message representation aggregation approach is introduced to aggregate the representations extracted from message pairs with high confidence levels, serving as the final representation of the message. This approach also avoids the over-smoothing of representations and enhances the similarity between representations within the same event. Finally, we design a new clustering constraint encompassing inter-cluster and intra-cluster losses. The inter-cluster loss aims to distance the cluster centers of different events, whereas the intra-cluster loss ensures that messages of the same event are gathered close to their respective cluster center. This clustering constraint strengthens the cohesiveness within clusters and the separation between them, thereby effectively improving the distinguishability of the message representations. By leveraging the inherent abilities of PLMs to learn, retain, and expand knowledge, we achieve incremental social event detection.
We conduct comprehensive experiments on three datasets, Events2012 (McMinn et al., 2013), Events2018 (Mazoyer et al., 2020), and Arabic-Twitter (Alharbi and Lee, 2021). Additionally, analyses and comparisons are conducted with one community detection method, seven GNN-based methods, and seven pre-training-based methods across various detection scenarios. Specifically, we first conduct comparisons and analyses with five mainstream social event detection methods: EventX (Liu et al., 2020a), PCGNN (Peng et al., 2019), KPGNN(Cao et al., 2021), QSGNN (Ren et al., 2022a), and FinEvent (Peng et al., 2022), in both offline and online scenarios, to demonstrate the effectiveness of the proposed $\mathrm{RPLM}_{SED}$. The results indicate that $\mathrm{RPLM}_{SED}$ surpasses current mainstream social event detection methodologies in all evaluation metrics. The visualization analysis and results further demonstrate that the message representations obtained through $\mathrm{RPLM}_{SED}$ are more explicit. Secondly, we explore the performance of $\mathrm{RPLM}_{SED}$ in low-resource language situations on the Arabic-Twitter dataset and additionally introduce a method proposed for low-resource scenarios, CLKD (Ren et al., 2021), as a comparative baseline. Compared to all benchmark models, $\mathrm{RPLM}_{SED}$ demonstrates superior model performance, highlighting its adaptability in low-resource environments. Furthermore, in the experiments on the long-tail distribution of social events, we additionally included three comparison baselines, MVGAN (Cui et al., 2021), ETGNN (Ren et al., 2022b), and $\mathrm{UCL}_{\mathrm{SED}}$ (Ren et al., 2023), for comparison. Compared to the currently best-performing models designed to tackle the long-tail distribution problem, $\mathrm{RPLM}_{SED}$ still demonstrated state-of-the-art (SOTA) performance. This further substantiates the efficacy of $\mathrm{RPLM}_{SED}$ in effectively detecting events in various scenarios. To investigate the ability of PLMs in the social event detection task, we undertook a comparative analysis of the performance of eight models in the online scenario: the Word2Vec (Mikolov et al., 2013) word embedding technique; BERT (Kenton and Toutanova, 2019) and RoBERTa (Liu et al., 2019); BERT-FT and RoBERTa-FT, which are fine-tuned versions of BERT and RoBERTa, respectively; two variants of $\mathrm{RPLM}_{SED}$, namely BERT-LR and RoBERTa-LR; and $\mathrm{RPLM}_{SED}$ itself. $\mathrm{RPLM}_{SED}$ consistently outperformed all comparison baselines. We also evaluate $\mathrm{RPLM}_{SED}$’s performance based on different PLMs, with results validating the effectiveness of leveraging PLMs for learning message representations. Additionally, we confirm each component’s critical role and indispensability within the $\mathrm{RPLM}_{SED}$ model through a series of ablation studies. Finally, by analyzing specific message block results, we validate the feasibility and stability of the $\mathrm{RPLM}_{SED}$.
@@ 2. Background and Overview
$S$
A social stream
$M$; $m$
A message block; A message in a message block or event
$E_{i}$
Set of events from message block $M_{i}$
$e_{k}$
@@
$\mathcal{G}$
A multi-relational message graph
$\mathcal{M}$
The set of all nodes in $\mathcal{G}$
$w$
The window size for maintaining the model
$\theta$; $\theta^{plm}$
The parameters of model; The parameters with the PLM section of model
$\zeta$
The parameter update weight during model maintenance
$\mathcal{R}_{t}$
Edge set with relational type $t$ in $\mathcal{G}$
$r_{i,s}^{t}$
An edge between $m_{i}$ and sampled node $m_{s}$ with relational type $t$
$r_{i,s}$
The multi-relational sequence between $m_{i}$ and sampled node $m_{s}$
$D$; $d$
Preprocessed message blocks; A message pair
$o$
The number of heads in structured attention mechanism
$\hat{x}_{i}$; $\hat{x}_{s}$
The representations of $m_{i}$ and sampled node $m_{s}$ obtained through structured attention mechanism and averaging
$q_{i}$
The similarity between $m_{i}$ and sampled node $m_{s}$
$X_{i}$; $Q_{i}$
A set of candidate representations of $m_{i}$; A set of similarities corresponding to $X_{i}$
$I_{i}$
A set of indices satisfying the similarity score
$\alpha$
The similarity threshold
$\overline{x}_{i}$
The final message representation of $m_{i}$
$B$
The training batch size
$H$; $h$
The center feature representation matrix; The central feature representation
$L_{pair}$
The pairwise cross-entropy loss
$L_{intra}$
The intra-cluster loss
$L_{inter}$
The inter-cluster loss
$L$
Total loss
$\kappa$; $\lambda$; $\mu$
Similarity pairwise cross-Entropy loss weight; Intra-cluster loss weight; Inter-cluster loss weight
@@ 2.2. Problem Definition
Figure 2. Incremental life-cycle in HP-Event architecture.
Figure 3. The architecture of the $\mathrm{RPLM}_{SED}$.
@@ 3.2. Pairwise Message Modeling Strategy
To jointly model social messages’ textual content and structural information, we model social messages into message pairs with multi-relational sequences. Firstly, we construct a Heterogeneous Information Network (HIN) (Peng et al., 2019, 2022), where each social message is treated as a message node. Concurrently, we extract a set of named entities, hashtags, and users (including users and mentioned users) as relational nodes from each message. Edges between the message nodes and the corresponding relational nodes denote their associations. For instance, the message $m_{1}$ depicted in the ‘Raw Messages’ part of Figure 3 includes three entities (‘Nobel’, ‘Mo Yan’, ‘Chinese’), one user (‘user1’), and one hashtag (‘Nobel’). Edges are established between $m_{1}$ and its related nodes. By merging duplicate nodes, a HIN graph is formed. In this graph, message nodes, entity nodes, user nodes, and hashtag nodes are represented by $m$, $e$, $u$, and $h$, respectively. Subsequently, the relational nodes in the heterogeneous information graph are converted into different types of edges between message nodes, thereby constructing a multi-relational message graph that includes single-type nodes and multiple-type relational edges. Additionally, edges are added between message nodes published within a 4-hour timeframe to capture temporal correlations. Finally, after establishing the multi-relational message graph, we further implement pairwise modeling for message nodes within the graph. For any given message node $m_{i}$, we sample $y$ positive and negative samples to form message pairs in the training phase and sample $n$ nodes in the detection phase. In this process, we prefer to select message nodes of the same event with richer connections to $m_{i}$ as positive sampling message nodes, whereas messages from different events with sparser connections to $m_{i}$ are chosen as negative sampling message nodes. This balanced and targeted sampling strategy helps to obtain more valuable sample pairs, enabling the model to learn more effective information during training. Considering that the multiple relations between messages may have varying impacts. Therefore, we comprehensively preserve these multifaceted relations through a multi-relational sequence. We define an external vocabulary to represent multi-relational edges ${\{\mathcal{R}_{t}\}\mid}_{t=1}^{T}$ in a multi-relational message graph. Each type of relation/edge $R_{t}$ occupies two distinct discrete values, indicating whether this relation/edge exists between messages. As illustrated in the ‘Pairwise Message Learning’ part of Figure 3, given a social message pair $(m_{i},m_{s})$, we map the edge relations between them into the corresponding multi-relational sequence as follows:
@@
Input: A social stream $S=M_{0},M_{1},M_{2},\cdots$, ground truth label set: $L=L_{0},L_{1},L_{2},\cdots$, training batch size $B$, number of iterations: $Epoch$, window size: $w$.
Output: Sets of events: $E_{0},E_{1},E_{2},\cdots$.
1 for $t=0,1,2,\cdots$ do
2       if $t\%w==0$ then
3             if $t==0$ then
4                   $M=M_{0}$;
5            else
6                   /* Parameter weighted update of PLM in $\mathrm{RPLM}_{SED}$. */
7                   $\theta_{t}^{PLM}=\zeta\cdot\theta_{t-w}^{PLM}+(1-\zeta)\cdot\theta_{0}^{PLM}$;
9                   $M=M_{t-w+1}+\cdots+M_{t}$;
10             end if
11            $Train=True$;
12             Execute Algorithm 2;
13             /* Model training */
14             for i in range(Epoch) do
15                   Execute Algorithm 3;
16
17             end for
18
19       end if
20      if $t~{}!=~{}0$ then
21            /* Detect events from $M_{t}$/
22             $M=M_{t},Train=False,L^{true}=L_{t}$;
23             Execute Algorithm 2;
24             Execute Algorithm 3;
25       end if
26
27 end for
Input: Message block: $M$, whether message block is a training dataset: $Train$.
Output: Train data blocks: $D^{t},D^{v}$ or detection data block: $D^{e}$.
1 if $Train==True$ then
2       $M^{t},M^{v}\leftarrow$ Split dataset proportionally from $M$;
3else
4       $M^{e}=M$;
5
6 end if
[…]
8 for $m~{}in~{}M^{t}$ do
9       for $a=1,2,3,\cdots,y$ do
10             $m^{+}_{s},m^{-}_{s}\leftarrow\{m,\text{Sample}(\mathcal{G}^{t}[\mathcal{M}])\}$;
11             $d^{+}$, $d^{-}\leftarrow$ Eq. 1, Eq. 2;
12             Append $d^{+}$, $d^{-}$ to $D^{t}$ (Eq. 3);
13       end for
14
15 end for
16for $m~{}in~{}M^{v}(or~{}M^{e})$ do
17       for $b=1,2,3,\cdots,n$ do
18             $m_{s}\leftarrow\text{Sample}(\mathcal{G}^{v}[\mathcal{M}])$; // or $\mathcal{G}^{e}[\mathcal{M}]$
19             $d\leftarrow$ Eq. 1, Eq. 2;
20             Append $d$ to $D^{v}$ ($D^{e}$) ;
21       end for
22
23 end for
return $D^{t}$, $D^{v}$, $D^{e}$;
Algorithm 2 Pairwise Message Modeling Strategy.

### v2

@@ Abstract.
Social Event Detection (SED) aims to identify significant events from social streams, and has a wide application ranging from public opinion analysis to risk management. In recent years, Graph Neural Network (GNN) based solutions have achieved state-of-the-art performance. However, GNN-based methods often struggle with missing and noisy edges between messages, affecting the quality of learned message embedding. Moreover, these methods statically initialize node embedding before training, which, in turn, limits the ability to learn from message texts and relations simultaneously. In this paper, we approach social event detection from a new perspective based on Pre-trained Language Models (PLMs), and present $\mathrm{RPLM}_{SED}$ (Relational prompt-based Pre-trained Language Models for Social Event Detection). We first propose a new pairwise message modeling strategy to construct social messages into message pairs with multi-relational sequences. Secondly, a new multi-relational prompt-based pairwise message learning mechanism is proposed to learn more comprehensive message representation from message pairs with multi-relational prompts using PLMs. Thirdly, we design a new clustering constraint to optimize the encoding process by enhancing intra-cluster compactness and inter-cluster dispersion, making the message representation more distinguishable. We evaluate the $\mathrm{RPLM}_{SED}$ on three real-world datasets, demonstrating that the $\mathrm{RPLM}_{SED}$ model achieves state-of-the-art performance in offline, online, low-resource, and long-tail distribution scenarios for social event detection tasks.
@@ 1. Introduction
However, GNN-based social event detection methods still exhibit the following limitations.
  1. (1)
    Graphs constructed from common attributes (e.g., entities, hashtags, users, etc.) between messages face issues of missing and noisy edges (Liu et al., 2022). These issues are primarily manifested in the lack of connections between message nodes within the same event and the existence of connections between nodes from different events. Whether modeling social messages as HIN graph (Figure 1(a)), homogeneous graph (Figure 1(b)), or multi-relational graph (Figure 1(c)), none of these approaches can handle well above issues. This challenge is multifaceted. On the one hand, intra-event messages may have no connections. This impedes GNNs from capturing associations between these nodes, resulting in an incomplete representation of related information. On the other hand, attribute co-occurrence is probable for inter-event messages, potentially introducing noise and impacting the quality of node representation. Besides, GNN-based methods fail to transmit information adequately in sparse graph scenarios, leading to suboptimal message representations.
  2. (2)
    GNN-based methods initialize node embeddings statically via pre-trained word embedding models (Mikolov et al., 2013; Pennington et al., 2014; Bojanowski et al., 2017) before training. This practice prevents the acquisition of better semantic embeddings from the message content in subsequent learning processes (Yang et al., 2021). Meanwhile, GNNs inherently separate the utilization of relations/edges and message content when learning message embeddings from explicit message graphs. The role of edges primarily select nodes for aggregation rather than directly participating in the processing of message representations. This approach may not fully consider the complexity of relations between messages and the potential interactions between these relations and message content (Ding et al., 2023).
  3. (3)
    GNN-based methods exhibit insufficient constraints on the training process, treating messages merely as targets to be pulled closer or pushed apart (Hermans et al., 2017).
@@
Herein, to circumvent the aforementioned challenges of GNN-based methods, we put forward a perspective of learning expressive message representations with Pre-trained Language Models (PLMs), instead of GNNs. Thus, we present a novel Relational prompt-based Pre-trained Language Model for Social Event Detection, namely $\mathrm{RPLM}_{SED}$. Concretely, we first propose a new pairwise message modeling strategy to model social messages. As illustrated in Figure 1(d), we sample other messages for each message to form pairs and convert the multiple relations between the two into a multi-relational sequence to preserve structural information. For any two messages, each type of relation between them and the existence or non-existence of these relations can be comprehensively represented as a multi-relational sequence. This strategy addresses the issues of missing and noisy edges in the message graph and achieves the joint modeling of the messages’ structure and semantics. Secondly, a new multi-relational prompt-based pairwise message learning mechanism is proposed to learn more detailed and comprehensive message representations from semantic and structural information between messages. It breaks away from traditional methods that rely on explicit relational graph learning and aggregation of message representations, and dividing message representation learning and aggregation into two independent parts. We transform the multi-relational sequences into multi-relational prompt embeddings and integrate them with the content embeddings of messages. By consistently encoding the embeddings of message pairs along with their corresponding multi-relational prompt embeddings through PLMs (Kenton and Toutanova, 2019; Liu et al., 2019; Zhang et al., 2023; Loureiro et al., 2023), we achieve the simultaneous utilization of message content and structural information, as well as the dynamic embedding and encoding of message content. To obtain more robust and stable message representations from their corresponding message pairs, a similarity-based message representation aggregation approach is introduced to aggregate the representations extracted from message pairs with high confidence levels, serving as the final representation of the message. This approach also avoids the over-smoothing of representations and enhances the similarity between representations within the same event. Finally, we design a new clustering constraint encompassing inter-cluster and intra-cluster losses. Unlike previous optimization methods that solely consider messages as targets, the objective of the clustering constraint optimization is focused on event clusters. The inter-cluster loss aims to distance the cluster centers of different events, whereas the intra-cluster loss ensures that messages of the same event are gathered close to their respective cluster center. This clustering constraint strengthens the cohesiveness within clusters and the separation between them, thereby effectively improving the distinguishability of the message representations. By leveraging the inherent abilities of PLMs to learn, retain, and expand knowledge, we achieve incremental social event detection.
We conduct comprehensive experiments on three datasets, Events2012 (McMinn et al., 2013), Events2018 (Mazoyer et al., 2020), and Arabic-Twitter (Alharbi and Lee, 2021). Analyses and comparisons are conducted with one community detection method (Liu et al., 2020a), eight GNN-based methods (Peng et al., 2019; Cao et al., 2021; Ren et al., 2022a; Peng et al., 2022; Ren et al., 2024; Cui et al., 2021; Ren et al., 2022b, 2023), and seven pre-training-based methods (Mikolov et al., 2013; Kenton and Toutanova, 2019; Liu et al., 2019) across online, offline, low-resource, and long-tail distribution scenarios. In all scenarios, $\mathrm{RPLM}_{SED}$ demonstrates SOTA performance compared to all benchmark methods. Additionally, we compare and analyze the performance of $\mathrm{RPLM}_{SED}$ based on different PLMs and report the time consumption for each part of $\mathrm{RPLM}_{SED}$. Finally, we analyze the performance of three density-based clustering algorithms (Campello et al., 2013; Ester et al., 1996; Ankerst et al., 1999) on the representations obtained by $\mathrm{RPLM}_{SED}$. The codes for all baseline models and $\mathrm{RPLM}_{SED}$, along with all datasets, are publicly accessible on GitHub [^1] https://github.com/RingBDStack/RPLM_SED.
@@ 2. Background and Overview
$S$ A social stream
$M$; $m$ A message block; A message in a message block or event
$E_{i}$ Set of events from message block $M_{i}$
@@
$\mathcal{G}$ A multi-relational message graph
$\mathcal{M}$ The set of all nodes in $\mathcal{G}$
$w$ The window size for maintaining the model
$\theta$; $\theta^{plm}$ The parameters of model; The parameters with the PLM section of model
$\zeta$ The parameter update weight during model maintenance
$\mathcal{R}_{t}$ Edge set with relational type $t$ in $\mathcal{G}$
$r_{i,s}^{t}$ An edge between $m_{i}$ and sampled node $m_{s}$ with relational type $t$
$r_{i,s}$ The multi-relational sequence between $m_{i}$ and sampled node $m_{s}$
$D$; $d$ Preprocessed message blocks; A message pair
$o$ The number of heads in structured attention mechanism
$\hat{x}_{i}$; $\hat{x}_{s}$ The representations of $m_{i}$ and sampled node $m_{s}$ obtained through structured attention mechanism and averaging
$q_{i}$ The similarity between $m_{i}$ and sampled node $m_{s}$
$X_{i}$; $Q_{i}$ A set of candidate representations of $m_{i}$; A set of similarities corresponding to $X_{i}$
$I_{i}$ A set of indices satisfying the similarity score
$\alpha$ The similarity threshold
$\overline{x}_{i}$ The final message representation of $m_{i}$
$B$ The training batch size
$H$; $h$ The center feature representation matrix; The central feature representation
$L_{pair}$ The pairwise cross-entropy loss
$L_{intra}$ The intra-cluster loss
$L_{inter}$ The inter-cluster loss
$L$ Total loss
$\kappa$; $\lambda$; $\mu$ Similarity pairwise cross-Entropy loss weight; Intra-cluster loss weight; Inter-cluster loss weight
@@ 2.2. Problem Definition
Figure 2. Incremental life-cycle in HP-Event architecture. Stage I. The initial data block $M_{0}$ is utilized to train the initial model. Stage II. The model detects events within each message block of the window separately. Stage III. The model is maintained through the current window’s message blocks. Then, return to the detection phase to detect events of each message block in the next window.
@@ 3. Methodology
Figure 3. The architecture of the $\mathrm{RPLM}_{SED}$. Raw Message shows real messages from Twitter. Pairwise Message Modeling (Section 3.2) initially constructs a multi-relational message graph, then sample massage to form pairs and map relations to the multi-relational sequence. Pairwise Message Learning (Section 3.3) embeds and encodes message content and relational sequence, then extracts message representations from the encoding vector. the messages’ similarity is assessed by the classifier. In the detection phase, candidate representations of each message are selected to aggregate the final representations. Optimization Objective (Section 3.4) comprises pairwise cross-entropy, inter-cluster and intra-cluster losses, jointly guiding model training.
@@ 3.2. Pairwise Message Modeling Strategy
To jointly model social messages’ textual content and structural information, we model social messages into message pairs with multi-relational sequences. Firstly, we construct a Heterogeneous Information Network (HIN) (Peng et al., 2019, 2022), where each social message is treated as a message node. Concurrently, we extract a set of named entities, hashtags, and users (including users and mentioned users) as relational nodes from each message. Edges between the message nodes and the corresponding relational nodes denote their associations. For instance, the message $m_{1}$ depicted in the ‘Raw Messages’ part of Figure 3 includes three entities (‘Nobel’, ‘Mo Yan’, ‘Chinese’), one user (‘user1’), and one hashtag (‘Nobel’). Edges are established between $m_{1}$ and its related nodes. By merging duplicate nodes, a HIN graph is formed. In this graph, message nodes, entity nodes, user nodes, and hashtag nodes are represented by $m$, $e$, $u$, and $h$, respectively. Subsequently, the relational nodes in the heterogeneous information graph are converted into different types of edges between message nodes, thereby constructing a multi-relational message graph that includes single-type nodes and multiple-type relational edges. Additionally, edges are added between message nodes published within a 4-hour timeframe to capture temporal correlations. Finally, after establishing the multi-relational message graph, we further implement pairwise modeling for message nodes within the graph. For any given message node $m_{i}$, we sample $y$ positive and negative samples to form message pairs in the training phase and sample $n$ nodes in the detection phase, where $n$ is three times $y$. In the training stage, we prefer to select message nodes of the same event with richer connections to $m_{i}$ as positive sampling message nodes, whereas messages from different events with sparser connections to $m_{i}$ are chosen as negative sampling message nodes. This balanced and targeted sampling strategy helps to obtain more valuable sample pairs, enabling the model to learn more effective information during training. In the detection phase, nodes that have richer connections to $m_{i}$ are sampled to form message pairs. Considering that the multiple relations between messages may have varying impacts. These multiple relations are preserved comprehensively through a multi-relational sequence. We define an external vocabulary to represent multi-relational edges ${\{\mathcal{R}_{t}\}\mid}_{t=1}^{T}$ in a multi-relational message graph. Each type of relation/edge $R_{t}$ occupies two distinct discrete values, indicating whether this relation/edge exists between messages. Consequently, The different states (exist or not) of different relations between messages correspond to unique discrete values. As illustrated in the ‘Pairwise Message Learning’ part of Figure 3, given a social message pair $(m_{i},m_{s})$, we map the edge relations between them into the corresponding multi-relational sequence as follows: